# MiraGe: Editable 2D Images using Gaussian Splatting

### v1

@@ 3 MiraGe: Editable 2D Images using Gaussian Splatting
To address this challenge, we employ a novel approach utilizing two opposing cameras positioned along the $Y$ axis, symmetrically aligned around the origin and directed towards one another. The first camera is tasked with reconstructing the original image, while the second models the mirror reflection. The photograph is thus conceptualized as a translucent tracing paper sheet, embedded within the 3D spatial context. The reflection can be effectively represented by horizontally flipping the image, denoted as $\mathcal{M}(I)$. This mirror-camera setup enhances the fidelity of the generated reflections, providing a robust solution for accurately capturing visual elements.
The MiraGe is initialized according to equation Eqn. 8 and utilizes a cost function: $\mathcal{L}(I)+L(\mathcal{M}(I))$. We simultaneously model both the image and its mirrored reflection, as shown in the second row in Fig. 6.
@@
Editability
The ability to manipulate Gaussians based on their spatial positioning empowers MiraGe to effectively edit 2D images. When utilizing a mirror camera, the quality of the resulting images is sufficiently high, enabling the parameterization and animation of Gaussians to significantly reduce artifacts. Our findings demonstrate that our model facilitates the animation of both segmented objects and entire scenes. Users can create manual animation, or leverage automated processes using physics engines like Taichi_elements or Blender (8, 9). To incorporate MiraGe with the 2D physics engine, we use 2D-MiraGe (Fig 8).
@@ 4 Experiments
We split the experimental section of our paper into two main parts. First, we demonstrate that our approach achieves high-quality 2D reconstruction by comparing it with existing models. Second, we highlight the versatility of our method in image editing both full scenes (Fig. 10) and selected objects (Fig. 1, 7), presenting examples of user-driven modifications and demonstrations involving physical simulations (Fig. 8, 9).
Figure 14: MiraGe model allows modifications in 3D space, but the model is limited by 2D images, which was used in training. When we move some elements from the foreground, we are not able to see the background since the model only reconstructs objects. After modification, we can see artifacts on the hind paw of the presented animals.
Reconstruction quality
To assess image quality, we utilized two established metrics, i.e., PSNR and MS-SSIM (Wang et al., 2003). These metrics measure the distortion between the reconstructed images and the ground truth.
Our image reconstruction assessment utilizes two widely-recognized datasets. Specifically, we employ the Kodak dataset [^2] https://r0k.us/graphics/kodak/, which includes 24 images at a resolution of $768\times 512$, alongside the DIV2K validation set (Agustsson & Timofte, 2017), which involves 2×2$\times$2 × bicubic downscaling and comprises 100 images with sizes ranging from $408\times 1020$ to $1020\times 1020$. The dataset was selected to facilitate direct comparison with the work of GaussianImage.
As a baselines we use competitive INR methods GaussianImage (Zhang et al., 2024), SIREN (Sitzmann et al., 2020b), WIRE (Saragadam et al., 2023b), I-NGP (Müller et al., 2022), and NeuRBF (Chen et al., 2023).
In Tab. 1, we demonstrate the performance outcomes of different methods on the Kodak and DIV2K datasets. We see that our proposition outperforms the previous solutions on both datasets. The quality measured by both metrics shows significant improvement compared to all the previous approaches. Fig. 11 illustrates a general trend observed during training in contest of image reconstruction. The selection of hyperparameters, including the number of iterations, was inspired by the principles of 3DGS. We provide ablation studies and extensive numerical analyses in the appendix for further insights.
@@
Due to our particular focus on animation, we analyzed FPS trends to benchmark real-time performance. Fig. 12 shows that while our model introduces a higher number of parameters, leading to a decrease in FPS compared to GaussianImage, it maintains the ability to render animations in real time
Manual modification
@@
MiraGe allows for manual manipulation of 2D images. By leveraging GaMeS parameterization, each Gaussian component is represented as a triangle. Vertices can then be independently adjusted and moved within 3D space, enabling flexible image modification (Fig. 27).
@@
We demonstrate examples of modifications using datasets such as DIV2K, Kodak, and Animals [^3] https://www.kaggle.com/datasets/alessiocorrado99/animals10. Additionally, we generated our own 2D images using DALL-E 3 to illustrate the benefits of our method. We can obtain modifications of small details like changing fingers position (Fig. 1), human facial expressions (Fig. 13) or dog poses (Fig. 14). As MiraGe can trained in a 3D context, we can implement modifications in the third dimension to create the illusion of 3D transformation (Fig. 36).
@@
It is crucial to note that when we displace elements from the foreground, the background remains unseen because the model only reconstructs the objects. This is demonstrated in Fig. 14, where artifacts are apparent on the hind paw of the depicted animal.
Table 1: Quantitative comparison with various baselines in PSNR and MS-SSIM. MiraGe gives state of the art results. Model-$x$ denotes that the model was initialized with $x$ Gaussians.
Kodak dataset DIV2K dataset
PSNR $\uparrow$ MS-SSIM $\uparrow$ PSNR $\uparrow$ MS-SSIM $\uparrow$
WIRE (Saragadam et al., 2023a) 41.47 0.9939 35.64 0.9511
SIREN (Sitzmann et al., 2020a) 40.83 0.9960 39.08 0.9958
I-NGP (Müller et al., 2022) 43.88 0.9976 37.06 0.9950
NeuRBF (Chen et al., 2023) 43.78 0.9964 38.60 0.9913
3DGS (Kerbl et al., 2023) 43.69 0.9991 39.36 0.9979
GaussianImage-70k (Zhang et al., 2024) 44.08 0.9985 39.53 0.9975
GaussianImage-100k* (Zhang et al., 2024) 38.93 0.9948 41.48 0.9981
MiraGe-70k (our) 57.41 0.9998 53.22 0.9996
MiraGe-100k (our) 59.52 0.9999 54.54 0.9998
Physics application in MiraGe
@@
Using 2D-MiraGe we can express Gaussian components with a 2D point cloud. Therefore, we can use MPM (Hu et al., 2018) based physics engine implemented, for example, in Taichi__\__elements. This high-performance physics engine supports multiple materials, including elastic objects and sand. We use inspiration from GASP (Borycki et al., 2024) and train simulation on 2D points then use physical deformation on triangle soup. In Fig. 8, we present simulation results obtained using Taichi_elements. As we can see, we can add physical properties to 2D objects. On the other hand, using Amorphous-MiraGe or Graphite-MiraGe we can use Blender and modify directly parameterized flat 3D Gaussian (Fig. 9).
@@ 5 Conclusion
In this paper, we introduce MiraGe that uses flat 3D Gaussian components to model 2D images. MiraGe gives state-of-the-art reconstruction quality and simultaneously allows image manipulation. Furthermore, we can modify photos on a plane (Fig. 8) and in 3D space (Fig. 10, 9). In consequence, we obtain the illusion of 3D-based modifications. Furthermore, we can combine our solution with a physics engine to obtain realistic motion in the image. Conducted experiments show that MiraGe is applicable in many different scenarios and produces high-quality simulations.
Limitation
@@
It is crucial to note that the model is not generative, so improper adjustment of Gaussian positions can cause gaps in the image (e.g. a missing dog’s paw 14). While the model can produce realistic changes, significant modification may introduce visual artifact. Moreover, our model requires encoding more parameters than GaussianImage to achieve high-quality image reconstruction for animation. Addressing this trade-off will be a focus of our future work.
Social impact
The model can be applied to generate novel image transformations used to dataset augmentation, facial recognition (Sanil et al., 2023), distortion correction, and visualizing architectural designs (Gerstweiler et al., 2018), as shown in Fig. 10. In the field of medical imaging, it is utilized for refining anatomical models and enhancing the accuracy of surgical simulations and diagnostic tools (Lin et al., 2023). Additionally, in 2D computer games, the model facilitates more realistic animations by incorporating physics-based effects (Mohd et al., 2023).

### v2

@@ 3 MiraGe: Editable 2D Images using Gaussian Splatting
Mirror camera We employ a novel approach utilizing two opposing cameras positioned along the $Y$ axis, symmetrically aligned around the origin and directed towards one another. The first camera is tasked with reconstructing the original image, while the second models the mirror reflection. We introduced the mirror camera to ensure that Gaussians remain confined within a specific spatial region between the cameras, enhancing control and precision.
Figure 11: MiraGe model allows modifications in 3D space, but the model is limited by 2D images, which was used in training. When we move some elements from the foreground, we cannot see the background since the model only reconstructs objects. Next, we can use image Inpainting to fill the missing parts, allowing for more realistic modifications.
The reflection can be effectively represented by horizontally flipping the image, denoted as $\mathcal{M}(I)$. This mirror-camera setup enhances the fidelity of the generated reflections, providing a robust solution for accurately capturing visual elements. We consider the additional camera as a means of augmenting the dataset to improve the accuracy of the representation. The MiraGe is initialized according to equation Eqn. 2 and utilizes a cost function: $\mathcal{L}(I)+L(\mathcal{M}(I))$. We simultaneously model both the image and its mirrored reflection, as shown in the second row in Fig. 6. We provide numerical comparisons in the ablation study in the Appendix.
@@
Editability The ability to manipulate Gaussians based on their spatial positioning empowers MiraGe to effectively edit 2D images. When utilizing a mirror camera, the quality of the resulting images is sufficiently high, enabling the parameterization and animation of Gaussians to significantly reduce artifacts. Our findings demonstrate that our model facilitates the animation of both segmented objects and entire scenes. Users can create manual animation or leverage automated processes using physics engines like Taichi_elements or Blender (Fig. 3,7). To incorporate MiraGe with the 2D physics engine, we use 2D-MiraGe (see, Fig. 7). In Fig. 8, we demonstrate that our method can also be applied to edit more complex scenes, such as changing human expression.
@@ 4 Experiments
We split the experimental section of our paper into two main parts. First, we demonstrate that our approach achieves high-quality 2D reconstruction by comparing it with existing models. Second, we highlight the versatility of our method in image editing both full scenes (Fig. 8) and selected objects (Fig. 1), presenting examples of user-driven modifications and demonstrations involving physical simulations (Fig. 3, 7).
Reconstruction quality Our image reconstruction assessment utilizes two widely recognized datasets. Specifically, we employ the Kodak dataset [^3] https://r0k.us/graphics/kodak/, which includes 24 images at a resolution of $768\times 512$, alongside the DIV2K validation set (Agustsson & Timofte, 2017), which involves 2×2$\times$2 × bicubic downscaling and comprises 100 images with sizes ranging from $408\times 1020$ to $1020\times 1020$. The dataset was selected to facilitate direct comparison with the work of GaussianImage. As a baselines we use competitive INR methods GaussianImage (Zhang et al., 2024), SIREN (Sitzmann et al., 2020b), WIRE (Saragadam et al., 2023b), I-NGP (Müller et al., 2022), and NeuRBF (Chen et al., 2023).
In Tab. 1, we demonstrate the performance outcomes of different methods on the Kodak and DIV2K datasets. We see that our proposition outperforms the previous solutions on both datasets. The quality measured by both metrics shows significant improvement compared to all the previous approaches. Fig. 9 illustrates a general trend observed during training in the contest of image reconstruction. The selection of hyperparameters, including the number of iterations, was inspired by the principles of 3DGS. We provide ablation studies and extensive numerical analyses in the appendix for further insights.
@@
Manual modification MiraGe allows for manual manipulation of 2D images. By leveraging GaMeS parameterization, each Gaussian component is represented as a triangle. Vertices can be independently adjusted and moved within 3D space, enabling flexible image modification (Fig. 2).
@@
We demonstrate examples of modifications using datasets such as DIV2K, Kodak, and Animals [^4] https://www.kaggle.com/datasets/alessiocorrado99/animals10. Additionally, we generated our own 2D images using DALL-E 3 to illustrate the benefits of our method. We can obtain modifications of small details like changing fingers’ position (Fig. 1), human facial expressions (Fig. 17) or dog poses (Fig. 11). As MiraGe can trained in a 3D context, we can implement modifications in the third dimension to create the illusion of 3D transformation (Fig. 36).
Figure 12: Comparison between PhysGen (Liu et al., 2025) and MiraGe. In the animation of a red block, reasonable doubts related to the correctness of the simulated physics of PhysGen arise when a reader follows the behavior of a red element upon hitting the wall; contrary to everyday experience, the front part (instead of the back part) of the figure bounces off the tabletop. In another animation, domino pieces start overlapping throughout the simulation. Our renders have properly solved the mentioned issues.
@@
It is crucial to note that when we displace elements from the foreground, the background remains unseen because the model only reconstructs the objects. This is demonstrated in Fig. 11, where artifacts are apparent on the hind paw of the animal depicted. To reduce such a problem, we can use Inpainting (Perche-Mahlow et al., 2024) on the image background.
We conducted a comparative analysis of our editing approach against the DragGAN model (Pan et al., 2023). Here, we focus on the ability to perform localized edits, such as closing the mouth, while preserving other features, such as dimples (see Fig. 8). The visual results, presented in Fig. 10, highlight key distinctions between the two models. As DragGAN is a generative model, modifications often result in unintended global transformations, for instance, attempting to adjust a leg’s position may inadvertently modify facial features. In contrast, our method demonstrates the ability to move elements like the leg with realistic results and without compromising other aspects of the image.
@@
Physics application in MiraGe Using 2D-MiraGe we can express Gaussian components with a 2D point cloud. Therefore, we can use an MPM (Hu et al., 2018) based physics engine implemented, for example, in Taichi__\__elements. This high-performance physics engine supports multiple materials. We use inspiration from GASP (Borycki et al., 2024) and train simulation on 2D points, then use physical deformation on triangle soup. In Fig. 7, we present simulation results obtained using Taichi_elements. As we can see, we can add physical properties to 2D objects. On the other hand, using Amorphous-MiraGe or Graphite-MiraGe, we can use Blender and modify directly parameterized flat 3D Gaussian (Fig. 3). Moreover, we compare MiraGe with PhysGen (Liu et al., 2025) (Fig. 12). Our method demonstrates better accuracy, ensuring that objects do not overlap. A key advantage is that users can directly and intuitively influence modifications.
@@ 5 Conclusion
In this paper, we introduce MiraGe that uses flat 3D Gaussian components to model 2D images. MiraGe gives state-of-the-art reconstruction quality and simultaneously allows image manipulation. Furthermore, we can modify photos on a plane (Fig. 7) and in 3D space (Fig. 3). In consequence, we obtain the illusion of 3D-based modifications. Furthermore, we can combine our solution with a physics engine to obtain realistic motion in the image. Conducted experiments show that MiraGe is applicable in many different scenarios and produces high-quality simulations.
@@
Limitation It is crucial to note that the model is not generative, so improper adjustment of Gaussian positions can cause gaps in the image (e.g. missing dog’s paw). This can be alleviated by using image Inpainting (Fig. 11). Although the model can produce realistic changes, a significant modification may introduce a visual artifact. Moreover, our model requires encoding more parameters than GaussianImage to achieve high-quality image reconstruction for animation. Addressing this trade-off will be a focus of our future work.
Appendix A appendix
Here, we provide a social impact and a comprehensive overview of the implementation details. Furthermore, we present supplementary experimental results, such as extended performance evaluations and ablation studies focusing on camera settings.
A.1 Social impact
Editing images in pixel representation is a well-established technique with many existing solutions, and 2D image editing tools (e.g. Photoshop, GIMP). MiraGe presents a concept that combines 2D and 3D representations to achieve the 2.5D effect commonly used in video games and VR (Feyer et al., 2024). The approach streamlines image reconstruction and manipulation without complex 3D generative models. It democratizes 3D editing, benefiting digital art, AR/VR and scientific visualization. By making 3D effects accessible and computationally efficient.
A.2 Implementation details
The source code for this project will be made publicly available on GitHub. Our code was developed based on the GaMeS framework (see Fig. 13) and is distributed under the GS Vanilla license. Computational experiments in the main paper were conducted using NVIDIA GeForce RTX 4070 Laptop version and NVIDIA GeForce RTX 2080. Appendix time comparisons were reported using NVIDIA GeForce RTX 2080.
Figure 15: Example of artifacts generated during animation, typically due to imperfect rendering. The model was trained on a white background, leaving residual white Gaussians along the border of the camel’s muzzle, leading to artifacts. In this instance, the Graphine-MiraGe performed best in handling the head-turning movement
Building upon the GaMeS framework, we initialized the Gaussian distributions to lie perpendicular to the $XZ$ plane. In our model, where all Gaussians are constrained to a 2D plane at rendering time, we consider only the rotation angle, denoted as $\phi$, as the primary rotation parameter. To facilitate the rendering of Gaussians positioned on the $XZ$ plane, $\phi$ serves as the primary learning parameter. The corresponding quaternions of rotation are computed as follows: for rotation about the x-axis $q_{x}=[cos(\frac{\phi}{2}),sin(\frac{\phi}{2}),0,0]$, and for the z-axis $q_{z}=[cos(\frac{\pi}{2}),0,0,sin(\frac{\pi}{2})]$. Since no rotation occurs about the y-axis, the quaternion remains $q_{y}=[1,0,0,0]$. These quaternions are then combined through multiplication to form a new rotation matrix, ensuring precise alignment of the Gaussians on the $XZ$ plane.
A.3 Supplementary Numerical Findings from the Primary Paper
We conducted an extensive analysis of the MiraGe model due to its unique ability to control the behavior of Gaussians. Three distinct settings for Gaussian movement were explored:
  • Amorphous the first allows Gaussians to move freely in 3D space,
  • 2D: the second restricts their movement to align parallel to the $XZ$ plane
  • Graphite the third confines all Gaussians to the $XZ$ plane, effectively creating a 3D representation.
Figure 18: MiraGe allows for manual image edits and for using a physics engine for real-life-like image modifications. The left image illustrates a Gaussian representation achieved through a triangle mesh triangle soup, while the accompanying point-based depiction provides finer details, offering a more refined visual comparison.
Figure 19: MiraGe allows for the modification of larger scenes. We can selectively alter specific areas and introduce smooth movements or material adjustments. In this example, the bottom of the blanket is shown in motion. This, along with other modifications, is available in the supplementary files as videos.
Table 2: Ablation study of the effect of adding the mirror camera as augmentation technique on training time and the output image quality measured in widely recognized metrics: PSNR, MS-SSIM, LPSIS. The experiment was performed with an initial 100k Gaussians.
Kodak dataset
Gaussian control method Camera Setting PSNR $\uparrow$ MS-SSIM $\uparrow$ LPSIS $\downarrow$ Training Time(s) $\downarrow$
Amorphous One camera 51.56 0.9996 0.0050 448.73
Mirror cameras 59.52 0.9999 0.0005 639.66
Graphite One camera 42.49 0.9948 0.2984 398.54
Mirror cameras 46.90 0.9983 0.1238 739.66
2D One camera 42.75 0.9950 0.2931 552.80
Mirror cameras 48.82 0.9987 0.0071 942.78
DIV2K dataset
Gaussian control method Camera Setting PSNR $\uparrow$ MS-SSIM $\uparrow$ LPSIS $\downarrow$ Training Time(s) $\downarrow$
Amorphous One camera 46.00 0.9991 0.0162 690.98
Mirror cameras 54.54 0.9998 0.0033 946.35
Graphite One camera 40.02 0.9949 0.0312 582.50
Mirror cameras 46.52 0.9986 0.0117 1082.41
2D One camera 39.99 0.9949 0.0310 869.62
Mirror cameras 46.32 0.9985 0.0124 1278.33
A qualitative analysis was performed, considering the impact of the mirror camera (see Tab. 2), as well as the effect of varying the number of initial Gaussians on the overall model behavior (see Tab. 3). We also examined the impact of the camera using the Frames Per Second (FPS) metric and storage memory (see Tab. 4). Given the ongoing development of various 3D Gaussian Splatting compression techniques, we employed the .spz [^5] https://github.com/nianticlabs/spz tool to effectively compress the data.
Due to our particular focus on animation, we analyzed FPS trends to benchmark real-time performance. Fig. 14 shows that while our model introduces a higher number of parameters, leading to a decrease in FPS compared to GaussianImage, it maintains the ability to render animations in real-time.
Figure 21: MiraGe simplifies intuitive editing of images, allowing transformations such as adjusting the tilt of a hand with minimal complexity. This is achieved by modifying the object along the third dimension.
Tab. 2 shows the mirror camera view as the augmentation technique significantly improves the representation’s fidelity of every proposed Gaussian method. This behavior can be detected with the help of any of the measured metrics, i.e., PSNR, MS-SSIM and LPSIS. The drawback of improving the image quality is that a longer training time is required. The ablation study presented in Tab. 3 similarly suggests that our model scales well with the number of Gaussians used during model initialization. The striking example here is an average 62.12 PSNR score achieved by the Amorphous method on the Kodak dataset. The price paid in time of training grows here slower, i.e., increasing the number of starting Gaussians by an order of magnitude results in more extended though comparable training period length.
A.4 Extension of examples modification and Artifacts
Animating a full scene can be non-trivial, but it is possible. Fig. 17 demonstrates how a painting can be enlarged to visualize the impact of its placement in a room, offering a clear view of the potential arrangement. It is also possible to animate small, localized areas of the image, as demonstrated in Fig. 17. For the facial animation, we utilized the Lattice modifier in Blender. MiraGe enables manual image editing and incorporates a physics engine for image modifications (Fig. 18, 19, 20). It is crucial to remember that if certain Gaussians are shifted without considering their dependencies on others, the image will be disrupted. Therefore, the relationships between the Gaussians must be carefully modeled. We demonstrate this concept with the example of children playing with a blanket (Fig. 19). Despite the movement of the blanket (as seen in the supplementary video), the image remains uninterrupted and coherent.
A simple editing concept using 3D is shown in Fig 21. Fig. 20 illustrates a sculpture where the movement of the hand is achieved by adjusting the position of the shield behind the warrior. The image representation, based on parameterized Gaussians, facilitates precise editing of fine details within the 3D space.
Integrating the representation into Blender can introduce automatic adjustments that may result in visual artifacts (Fig. 15), particularly when training on images with a white background. These modifications can lead to unrealistic renderings that are challenging to detect through automated means and currently require subjective evaluation by a human observer.
Table 3: Measuring the influence of the initial number of Gaussians on the image reconstruction quality. The experiment was performed using a mirror camera view for every table entry.
Kodak dataset
Gaussian control method Initial Gaussians PSNR $\uparrow$ MS-SSIM $\uparrow$ LPSIS $\downarrow$ Training Time(s) $\downarrow$
Amorphous 10k 50.66 0.9987 0.3531 584.84
50k 55.54 0.9997 0.0033 634.65
100k 59.52 0.9999 0.0005 639.66
150k 62.12 0.9999 0.0002 676.10
Graphite 10k 40.39 0.9940 0.0599 651.32
50k 44.90 0.9973 0.2024 732.91
100k 46.90 0.9983 0.1238 739.66
150k 48.16 0.9987 0.0105 801.18
2D 10k 39.75 0.9886 0.0769 857.30
50k 45.03 0.9955 0.2789 876.56
100k 48.82 0.9987 0.0071 942.78
150k 50.54 0.9992 0.0031 955.86
DIV2K dataset
Gaussian control method Initial Gaussians PSNR $\uparrow$ MS-SSIM $\uparrow$ LPSIS $\downarrow$ Training Time(s) $\downarrow$
Amorphous 10k 49.53 0.9987 0.0322 852.19
50k 52.23 0.9995 0.0112 902.80
100k 54.54 0.9998 0.0033 946.35
150k 56.40 0.9999 0.0014 975.44
Graphite 10k 40.75 0.9959 0.0457 983.41
50k 44.67 0.9980 0.0216 1008.52
100k 46.52 0.9986 0.0117 1082.41
150k 47.61 0.9989 0.0083 1103.69
2D 10k 38.40 0.9920 0.0616 1166.09
50k 42.86 0.9967 0.0275 1256.62
100k 46.32 0.9985 0.0124 1278.33
150k 48.46 0.9990 0.0065 1415.54
Table 4: Ablation study of the effect of adding the mirror camera as augmentation technique on Kodak dataset measured using Frames Per Second (FPS) and memory storage.
Kodak dataset
Gaussian control method Camera Setting FPS Memory (MB) Compressed memory (MB)
Amorphous One camera 583.28 31.25 2.42
Mirror cameras 620.10 117.25 7.80
Graphite One camera 1157.75 30.71 2.68
Mirror cameras 650.75 117.91 9.22
2D One camera 1130.08 30.69 2.68
Mirror cameras 418.39 173.64 12.82